OpenStudy (anonymous):

Write the following statement as an equation and solve. Jake Drewrey, owner of Jake's Landscape Service, knows that one of his fertilizer tanks holds four times as many gallons of liquid fertilizer as a second tank. The two tanks together hold 362 gal. If both tanks are filled to capacity, how many gallons of fertilizer does each tank hold?

OpenStudy (texaschic101):

1st tank = x 2nd tank = y x + y = 362 x = 4y 4y + y = 362 5y = 362 y = 72.4 gallons x + y = 362 x + 72.4 = 362 x = 362 - 72.4 x = 289.6 gallons check.. x = 4y 289.6 = 4(72.4) 289.6 = 289.60 (correct) 1st tank (x) = 289.6 gallons 2nd tank (y) = 72.4 gallons

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

How do you write the statement as an equation?

OpenStudy (texaschic101):

I wrote it as 2 equations, then subbed, and solved.

OpenStudy (texaschic101):

I am not sure how to write it as 1 equation

OpenStudy (texaschic101):

@Loser66 ...can you take a peek at this

OpenStudy (texaschic101):

@SolomonZelman ...did I do this correctly ? Can it be written 1 equation ?

OpenStudy (texaschic101):

give them a minute...someone will come

OpenStudy (texaschic101):

@triciaal ...help

OpenStudy (texaschic101):

@perl...can you take a look at this please

OpenStudy (triciaal):

the above is correct@texaschic101 the second tank holds x x + 4x = 362 x = 362/5 = 72.4 gallons the first tank holds 4 * 72.4= same as (362.0 - 72.4) =

OpenStudy (triciaal):

@texaschic101